# UnsafeBench: Benchmarking Image Safety Classifiers on Real-World and AI-Generated Images

### v1

@@ Introduction
The general idea is to leverage the knowledge of pre-trained models, such as CLIP or VLMs, and adapt them to identify unsafe images under the guidance of annotated images. We elaborate on specific methodologies employed with each pre-trained model as follows.
Adapting CLIP for image classification tasks is a popular choice. Among the five conventional classifiers tested in UnsafeBench, four of them (Q16, MultiHeaded, NSFW_Detector, SD_Filter) utilize CLIP for image feature extraction. Similar to these classifiers, we employ linear probing and prompt learning to adapt CLIP to identify unsafe images. Linear probing [57] is a technique that adds a linear probing layer on top of a pre-trained model, allowing for fine-tuning on specific tasks, and prompt learning [34, 42] refers to a process that adapts a pre-trained model for various downstream tasks by optimizing prompts. For linear probing, we train 11 linear probing layers, each one identifying images from a specific unsafe category. Similarly, we optimize 11 sets of soft prompts using prompt learning. In the inference phase, according to the user’s defined scope of unsafe content, only the linear probing layers or soft prompts associated with the selected unsafe categories will be activated to determine if the image is safe or not.
Different from classifiers using linear probing and prompt learning, VLMs are generative models that output more information rather than simply logit probabilities. If we fine-tune the VLM to generate a response that indicates both the safety status and the specific unsafe category, we can build a unified model with VLMs like LLaVA instead of training classifiers separately for different categories. To this end, we employ supervised fine-tuning to adapt LLaVA for identifying unsafe images. We first build the fine-tuning dataset based on our annotated images. The dataset contains triplet elements: image, prompt, and target output. We adopt a similar prompt template as LLaMA-Guard [44] (an LLM-based safeguard), where we define the unsafe image taxonomy with specific unsafe categories and output format. Figure 9 demonstrates an example of the LLaVA fine-tuning dataset.
Since fine-tuning the entire LLaVA is computationally intensive, we resort to the Low-Rank Adaptation (LoRA) technique [43]. LoRA is a parameter-efficient fine-tuning technique that introduces a small set of additional trainable parameters to the model. It updates only these extra parameters while keeping the original model parameters frozen. With our fine-tuning dataset and the LoRA technique, we train a unified model that simultaneously predicts the safety status and identifies the specific unsafe category. The probability value is extracted from the logits associated with the next token prediction.
Figure 9: An example of LLaVA fine-tuning dataset. “<image>” is the visual token, representing the provided image. “Human” presents a prompt that describes the image safety classification task including the unsafe image taxonomy and the required output format. “Assistant” provides the target output.
@@
Datasets. We randomly split the UnsafeBench dataset into an 80% training and 20% testing ratio. To assess the model’s generalizability, we gather multiple external datasets for further evaluation: SMID [24], MultiHeaded Dataset [14], NudeNet Dataset [18], Hate Symbols [2], Self-Harm Dataset [22, 23], and Violent Behavior Dataset [28]. We outline these datasets and detailed experimental setup in Section A.2 and Section A.3 in the Appendix.
Evaluation on UnsafeBench Test Set. Table 6 shows the evaluation performance of PerspectiveVision models on the UnsafeBench test set. Among the three methods employed, linear probing (CLIP) and LoRA fine-tuning LLaVA demonstrate a higher fitting level on the UnsafeBench test set, with an overall F1-Score of 0.859 and 0.844. Meanwhile, prompt learning (CLIP) yields a lower F1-Score of 0.687. We observe that these models perform particularly well in identifying images within the Sexual, Political, Spam, and Shocking categories, where the F1-Scores are high, close to or exceeding 0.90.
Evaluation on External Datasets. Table 7 presents the evaluation results of the PerspectiveVision models across six external datasets. Here, we benchmark GPT-4V against these datasets to establish a high-performing baseline. GPT-4V demonstrates an overall F1-Score of 0.818, particularly standing out when classifying NudeNet images and Hate symbols. Prompt learning (CLIP) delivers comparable performance, with an overall F1-Score of 0.810. It shows a strong generalization capability in identifying morally negative, hate symbols, self-harm, and violent images. Following closely is the LoRA fine-tuned LLaVA, which achieves an F1-Score of 0.794. It surpasses other models in identifying AI-generated unsafe images from the MultiHeaded dataset. To conclude, prompt learning (CLIP) and LoRA fine-tuned LLaVA demonstrate better generalizability on external evaluation datasets, serving as the primary backbone models for PerspectiveVision.
Take-Aways. We build a comprehensive image classifier, PerspectiveVision, which identifies unsafe images across 11 unsafe categories. Similar to Google’s Perspective API, users can customize their definition of unsafe images by selecting from various categories. Based on our evaluation, PerspectiveVision performs comparably to zero-shot GPT-4V in identifying unsafe images, with its best backbone achieving an F1-Score of 0.810 on external evaluation datasets. We hope it serves as a baseline and will make it publicly available.
@@
Despite the variety of these solutions, employing open-source classifiers is a common practice within the research community to identify unsafe images. However, their performances on real-world images are underexplored, largely due to the absence of large labeled datasets. In our study, we first construct a comprehensive dataset encompassing a broad spectrum of unsafe content. We then thoroughly analyze their performances, including the effectiveness across different unsafe categories and the robustness against adversarial examples.
Counteracting AI-Generated Unsafe Images. Since text-to-image models like Stable Diffusion gained popularity in 2022, concerns have been raised regarding their risks of generating realistic unsafe images. Plenty of studies [62, 55, 73, 74, 30] focus on assessing these models’ risks. Schramowski et al. [62] take the first step in estimating the probability of Stable Diffusion in generating unsafe images when providing harmful prompts. Qu et al. [55] adopt a similar approach and find that text-to-image models are prone to generate sexually explicit, violent, disturbing, hateful, and political images. Other researchers investigate the proactive generation of unsafe images from text-to-image models through various attacks, such as data poisoning attacks [73] and adversarial examples [74, 30]. To mitigate the risks, another line of research focuses on enhancing safety measures. For example, Schramowski et al. [62] propose safe latent diffusion, which steers the generated images away from a list of unsafe concepts during the generation process. Guo et al. [41] takes the first step in using VLMs to identify unsafe images from user-generated content games. All the above works rely on existing image safety classifiers or VLMs to identify the unsafe images generated by text-to-image models. However, since these classifiers are mostly trained on real-world images, it is unclear how effectively they generalize to AI-generated images. Our benchmarking framework, UnsafeBench, investigates their ability to generalize to AI-generated unsafe images and explores the distinct characteristics of those images.
@@
We establish UnsafeBench, a benchmarking framework that comprehensively evaluates the performance of image safety classifiers on both real-world and AI-generated images. We construct a large image safety dataset of 10K manually annotated images and evaluate five conventional classifiers and three VLMs. Our analysis reveals that current image safety classifiers are constrained by imbalanced performance across different types of unsafe content and potential performance degradation on AI-generated images. We therefore introduce a comprehensive image moderation tool, PerspectiveVision, to evaluate the safety of a given image. Our work has the following implications for the research community, platform moderators, and model providers.
@@
Importance of UnsafeBench. One important contribution of UnsafeBench is the comprehensive dataset of 10K images, which are manually annotated as safe or unsafe. Besides these annotations, they can also be categorized into 11 unsafe categories or two sources (real-world or AI-generated). These attributes largely improve the dataset’s utility, enabling studying unsafe images in a multifaceted manner. We hope the dataset serves as a solid foundation for future image safety research. Another important contribution is the performance comparison among different image safety classifiers, which assists future studies in selecting the most suitable classifier according to the target unsafe content.
Enhancing Effectiveness on AI-Generated Images. Our analysis reveals that there exists a performance degradation when identifying unsafe AI-generated images, especially for classifiers trained merely on real-world images. This degradation is largely due to the distribution shift between two groups of images. Specifically, AI-generated images present distinct characteristics, such as artistic representations of unsafe content and grid layout. For platform moderators and model developers, it is necessary to account for AI-generated unsafe images in the era of generative AI. Extending the training dataset of image safety classifiers with AI-generated examples enhances generalizability. At the same time, we hope the UnsafeBench dataset aids future research to uncover more characteristics of AI-generated unsafe images.
Utility of PerspectiveVision. The results from UnsafeBench suggest that, currently, there is a lack of an open-source image safety classifier capable of effectively detecting a wide range of unsafe content. Inspired by Google’s Perspective API, which evaluates the toxicity of the given text from multiple perspectives, we establish a similar tool in the image safety domain, PerspectiveVision, as the first attempt. Similar to Perspective API, our tool enables users to customize the scope of unsafe content. The research community and platform moderators can directly apply it for various moderation tasks, such as reporting online unsafe images and evaluating the risks of new text-to-image models.
@@
Limitations. Our work has limitations. First, the UnsafeBench dataset is biased towards unsafe content, as it is curated by retrieving images most related to unsafe keywords. Second, images sourced from Lexica are mostly generated by a single text-to-image model, Stable Diffusion. We will expand our dataset with images generated by other text-to-image models and routinely update it. Third, we only evaluate the generalizability of PerspectiveVision on a limited range of unsafe categories due to the lack of labeled datasets. We will continue expanding the evaluation with more datasets and improving the model in future work.
@@ Appendix A Appendix
$\Delta x_{t}=\frac{f_{\theta}(x_{t})}{\|\nabla f_{\theta}(x_{t})\|_{2}^{2}}\nabla f_{\theta}(x_{t}).$ (5)
@@
Hate Symbols [2] are the images and signs that are commonly used by hate groups to spread hateful ideology. For example, Happy Merchant is an anti-semitic symbol that defames the Jewish community. We collect 1,084 hate symbols from the Anti-Defamation League (ADL) [2], an international anti-hate organization.
@@
  • Linear Probing. We independently train 11 MLPs, each designated to identify one of the 11 unsafe categories. We load the entire set of training images to train each MLP. In this case, images classified as unsafe in other categories are still treated as safe. This ensures that each MLP focuses exclusively on one unsafe category. During training, we use the standard CrossEntropy loss and the Adam optimizer with a learning rate of 2e-4. The batch size is set to 128, and we train each MLP for 30 epochs.
@@
  • LoRA Fine-Tuning. We build the instruction dataset using the prompt template shown in Figure 9. Specifically, we construct a pair of instructional data points for each unsafe image, which we call the positive example and the negative example. In the positive example, given the unsafe image, we first identify the unsafe category it violates and then randomly append K (from 1 to 10) other unsafe categories to create a customized image safety taxonomy. The safety label is still recorded as unsafe. Meanwhile, in the negative example, we randomly select K unsafe categories that the given unsafe image does not violate, and then we record the safety label as safe. This approach ensures that VLMs learn the safety of images rather than simply memorizing the output format. We then fine-tune LLaVA on this dataset with its original training loss (CrossEntropy loss) and optimizer (AdamW), following the setup in [48]. The batch size is 16, and the learning rate is 2e-4.
@@
(a) LAION-5B
(b) Lexica
Figure 10: Examples of unsafe images from LAION-5B and Lexica. We randomly select these examples from the UnsafeBench dataset and blur the sexual images for censoring purposes.

### v2

@@ Introduction
Figure 10: An example of LLaVA fine-tuning dataset.
The main idea is to build a unified model with VLMs like LLaVA instead of training classifiers separately for different categories. To this end, we fine-tune a VLM, i.e., LLaVA, to identify unsafe images by generating a response that indicates both the safety status and the specific unsafe category. Since fine-tuning the entire LLaVA is computationally intensive, we use Low-Rank Adaptation (LoRA) [40] while training.
Training Dataset Construction. The training dataset is based on our annotated images. The dataset contains triplet elements: image, prompt, and target output. We adopt a similar prompt template as LLaMA-Guard [41] (an LLM-based safeguard), where we define the unsafe image taxonomy with specific unsafe categories and output format. Figure 10 demonstrates an example of the LLaVA fine-tuning dataset. To reduce overfitting, for each annotated image, we randomly remove K irrelevant categories and shuffle the order of the rest categories to build the prompt.
Training Dataset Augmentation. To further improve generalizability, we augment the dataset by label flipping and class balancing. On top of the basic training dataset, for images that are annotated as unsafe, we randomly sample K irrelevant categories to serve as the unsafe image taxonomy in the prompt. We ensure that these images are not unsafe within these categories and then flip the target output to the safe class. This helps the model learn the association between user-defined taxonomy in the prompt and the image label. Finally, we balance the number of samples between two classes by supplementing the smaller class.
Baselines
We employ two common techniques as baselines: linear probing and prompt learning to adapt CLIP to identify unsafe images. Linear probing [54] adds a linear probing layer on top of CLIP, allowing for fine-tuning on specific tasks, and prompt learning [31, 39] refers to a process that adapts CLIP for various downstream tasks by optimizing prompts. For linear probing, we train 11 linear probing layers, each one identifying images from a specific unsafe category. Similarly, we optimize 11 sets of soft prompts using prompt learning.
@@
Evaluation Datasets. We randomly split the UnsafeBench dataset into an 80% training and 20% testing ratio and take the test split as in-distribution evaluation dataset. To assess the model’s generalizability, we further collect multiple external datasets mainly from the training data of conventional classifiers. We regard them as out-of-distribution datasets: SMID [22], MultiHeaded Dataset [12], NudeNet Dataset [16], Self-Harm Dataset [20, 21], and Violent Behavior Dataset [25]. We outline these datasets and detailed experimental setup in Section A.4 and Section A.5 in the Appendix.
Effectiveness Evaluation. Table 6 presents the evaluation results of the PerspectiveVision models across UnsafeBench test set and five OOD datasets. Here, we benchmark GPT-4V against these datasets to establish a high-performing baseline. Among the three methods employed, linear probing (CLIP) and LoRA fine-tuning LLaVA demonstrate a higher fitting level on the UnsafeBench test set, with an overall F1-Score of 0.859 and 0.844. Meanwhile, on external datasets, LoRA fine-tuned LLaVA achieves the highest average F1-Score, 0.836. Although closely followed by GPT-4V, which has an F1-Score of 0.830, our model is notable for its open-source nature, large-scale detection, and fine-grained classification of different unsafe categories. Particularly, it surpasses other models in identifying unsafe images from the MultiHeaded dataset, which contains 800 AI-generated images by different text-to-image models. This improvement can be attributed to the inclusion of a large number of AI-generated images from UnsafeBench in its training data.
Table 7: Robust accuracy of PerspectiveVision.
Dataset GN FGSM PGD DeepFool Average
UnsafeBench-Test 0.986 0.960 0.926 0.936 0.952
MultiHeaded-D 0.966 0.912 0.886 0.884 0.912
Robustness Evaluation. We also test the robust accuracy of LoRA fine-tuned LLaVA against Gaussian and adversarial perturbations. Following the same setup in robustness evaluation, we randomly extract 500 correctly classified images from the UnsafeBench test set and an OOD AI-generated dataset, MultiHeaded-D, and craft adversarial examples with the same perturbation budget. We find that LoRA fine-tuning significantly improves robustness in this image safety classification task. While zero-shot LLaVA shows an average RA of 0.666-0.725 (see Table 5) under four perturbations, fine-tuned LLaVA increases the average RA to 0.952 on UnsafeBench test set and 0.912 on OOD AI-generated images, in Table 7.
Take-Aways. We build a comprehensive image classifier, PerspectiveVision, which identifies unsafe images across 11 categories by fine-tuning LLaVA on our UnsafeBench dataset. The inclusion of diverse AI-generated images not only enhances PerspectiveVision’s effectiveness, particularly with AI-generated images, but also significantly improves its robustness under the same strength of adversarial attacks.
@@
Despite the variety of these solutions, employing open-source classifiers is a common practice within the research community to identify unsafe images. However, their performances on real-world images are under-explored, largely due to the absence of large labeled datasets. In our study, we first construct a comprehensive dataset encompassing a broad spectrum of unsafe content. We then thoroughly analyze their performances, including the effectiveness across different unsafe categories and the robustness against adversarial examples.
Counteracting AI-Generated Unsafe Images. Since text-to-image models like Stable Diffusion gained popularity in 2022, concerns have been raised regarding their risks of generating realistic unsafe images. Plenty of studies [59, 52, 70, 71, 27] focus on assessing these models’ risks. Schramowski et al. [59] take the first step in estimating the probability of Stable Diffusion in generating unsafe images when providing harmful prompts. Qu et al. [52] adopt a similar approach and find that text-to-image models are prone to generate sexually explicit, violent, disturbing, hateful, and political images. Other researchers investigate the proactive generation of unsafe images from text-to-image models through various attacks, such as data poisoning attacks [70] and adversarial examples [71, 27]. To mitigate the risks, another line of research focuses on enhancing safety measures. For example, Schramowski et al. [59] propose safe latent diffusion, which steers the generated images away from a list of unsafe concepts during the generation process. Guo et al. [38] takes the first step in using VLMs and chain-of-thought to identify unsafe images from user-generated content in games. All the above works rely on existing image safety classifiers or VLMs to identify the unsafe images generated by text-to-image models. However, since these classifiers are mostly trained on real-world images, it is unclear how effectively they generalize to AI-generated images. Our benchmarking framework, UnsafeBench, investigates their ability to generalize to AI-generated unsafe images and explores the distinct characteristics of those images.
@@
We establish UnsafeBench, a benchmarking framework that comprehensively evaluates the effectiveness and robustness of image safety classifiers on both real-world and AI-generated images. We construct a large image safety dataset of 10K manually annotated images and evaluate five conventional classifiers and three VLMs. Our analysis reveals the drawbacks of these classifiers and the unique challenges posed by AI-generated images to both their effectiveness and robustness. We therefore introduce an improved image moderation tool, PerspectiveVision, that accounts for both real-world and AI-generated images. Our work has the following implications for the research community and platform moderators.
@@
Importance of UnsafeBench Dataset. We provide the first large-scale image safety dataset with 11 categories of unsafe content and two sources. Annotated images in this dataset are equipped with three “labels”: 1) safe/unsafe, 2) specific category, and 3) sourced from the real world or AI. These attributes greatly improve the dataset’s utility, enabling future studies to investigate unsafe images in a multifaceted manner. We hope the dataset serves as a solid foundation for future image safety research.
Challenges Posed by AI-Generated Images. In the era of generative AI, the influence of AI-generated content like unsafe imagery and deepfakes can be complex and multifaceted to existing models. Our analysis reveals that they do not only reduce the effectiveness of these models but also render them more vulnerable to adversarial attacks than real-world images. This is because, although current AI-generated images are realistic, there remains a distribution shift from real-world images in image quality statistics (noise level, entropy, etc.), styles, and layouts. For platform moderators and model developers, it is essential to include AI-generated unsafe images in the training data to mitigate these negative impacts.
Utility of PerspectiveVision. PerspectiveVision is built to address the drawbacks of existing image safety classifiers, including imbalanced performance across categories, state-of-the-art model (GPT-4V) being closed-source, and degraded performances on AI-generated images. It is also the first attempt to build an image moderation tool that covers both the wide spectrum of unsafe content and unsafe images from different sources. The research community and platform moderators can directly apply it for various moderation tasks, such as reporting online unsafe images and evaluating the risks of new text-to-image models.
@@
Limitations. Our work has limitations. First, the UnsafeBench dataset is annotated by only three authors, with the final label determined by majority vote. We did not rely on crowdsourcing workers for two reasons: 1) annotation requires expert knowledge in the image safety domain, such as identifying hate symbols, which cannot be guaranteed and would require specific training for crowdsourcing workers; 2) due to ethical considerations, we aim to minimize the number of annotators exposed to such unsafe content and prevent them from exposing it to third parties. Second, images sourced from Lexica are mostly generated by a single text-to-image model, Stable Diffusion. We will expand our dataset with images generated by other text-to-image models and routinely update it. Third, we only evaluate the generalizability of PerspectiveVision on a limited range of unsafe categories due to the lack of labeled datasets. We will continue expanding the evaluation with more datasets and improving the model in future work.
@@ Appendix A Appendix
Details of Conventional Classifiers
Details of Large VLMs
@@
$\Delta x_{t}=\frac{f_{\theta}(x_{t})}{\|\nabla f_{\theta}(x_{t})\|_{\infty}}\operatorname{sign}(\nabla f_{\theta}(x_{t})).$ (5)
@@
  • Linear Probing. We independently train 11 MLPs, each designated to identify one of the 11 unsafe categories. We load the entire set of training images to train each MLP. In this case, images classified as unsafe in other categories are still treated as safe. This ensures that each MLP focuses exclusively on one unsafe category. During training, we use the standard cross-entropy loss and the Adam optimizer with a learning rate of 2e-4. The batch size is set to 128, and we train each MLP for 30 epochs.
@@
  • LoRA Fine-Tuning. We build the instruction dataset using the prompt template shown in Figure 10. Specifically, we construct a pair of instructional data points for each unsafe image, which we call the positive example and the negative example. In the positive example, given the unsafe image, we first identify the unsafe category it violates and then randomly append K (from 1 to 10) other unsafe categories to create a customized image safety taxonomy. The safety label is still recorded as unsafe. Meanwhile, in the negative example, we randomly select K unsafe categories that the given unsafe image does not violate, and then we record the safety label as safe. This approach ensures that VLMs learn the safety of images rather than simply memorizing the output format. We then fine-tune LLaVA on this dataset with its original training loss (CrossEntropy loss) and optimizer (AdamW), following the setup in [45]. The batch size is 16, and the learning rate is 2e-4. We set the training epoch to 1.
@@
Figure 13: Image quality statistics of real-world and AI-generated images. We report signal-to-noise ratio (SNR), edge density, and entropy.
@@
Figure 15: Loss and loss change for four conventional classifiers on original images and their adversarial examples (FGSM). Classifiers tend to show higher loss values and greater loss changes with AI-generated images than with real-world images.